An experiment to measure the enthalpy change for the reaction of aqueous copper(II) sulfate, CuSO4(aq) and zinc, Zn(s) was carried out in a coffee cup calorimeter as follows.
Cu 2+(aq) + Zn(s) → Cu(s) + Zn2+(aq) 50.0 cm3 of 1.00 mol dm–3 copper(II) sulfate solution was placed in a polystyrene cup and its average temperature after three readings was noted to be 25OC and the zinc powder was added . The final temperature was noted to be 91.5 OC. Calulate the heat of the reaction in kJ
1
Expert's answer
2021-09-07T23:28:31-0400

"\u2206T = 91.5 - 25 =66.5\u00b0C"



"Moles = 1.00×\frac{50.0}{1000}=0.05mole"


"Mass= 159.609\u00d70.05=7.98grams"



"Heat = 4.18 \u00d7 7.98 \u00d7 66.5 =2,218.2kJ"